# Quantum cryptography beyond key distribution: theory and experiment

### v1

@@ IX.2 Verifiable computation
Figure 8: Verifiable blind quantum computing. First proposed in Fitzsimons and Kashefi (2017) this figure was taken from the talk prepared for presenting Gheorghiu et al. (2019), and reproduced with permission from the authors. Red qubits denote the computation qubits, yellow qubits are dummy qubits that disentangle their neighbors from the graph and orange qubits are trap qubits. Trap qubits produce determinstic measurement outcomes, so the client can check if the received outcome $b_{i}$ matches the honest execution.
@@ IX.3 Computation with classical clients
Assuming two quantum servers that share EPR pairs but cannot communicate with one another, a blind implementation of Shor’s factoring algorithm was performed for the number $N=15$, instructed by a fully classical client Huang et al. (2017). This involves the client splitting the computation into two parts, delegating one part to each of the photonic servers. Note that the CNOT gate is implemented differently for each of the two photonic servers, whose roles in the computation are asymmetric: for one server, it suffices to probabilistically post-select on the outputs of a polarizing beamsplitter, while the other server follows an implementation closer to Kiesel et al. (2005). By randomly alternating between computation and dummy tasks, the classical client is able to verify the behaviour of the servers and performing a CHSH test based on the protocol from Reichardt et al. (2013). Following up on the proposal from Huang et al. (2017), an extension to the blind factorization of $N=21$ was proposed Das and Sanders (2022).
@@ X.1 Composability
Composable quantum cryptography is a large field and goes beyond the scope of this review [^16] Even the framework that needs to be used for composable security, suitable for quantum multiparty protocols, is an open area of research.. Most multiparty protocols cannot be composably secure in the plain model (without assumptions such as the CRS), and this persists in the quantum case. For instance, it was shown that weak coin flipping is not composably secure in the plain model Wu et al. (2024), while stand-alone sequential composition is proven secure. The case of relativistic quantum protocols (for bit commitment and coin flipping) were also shown to be non-composable in the plain model Vilasini et al. (2019). Introducing assumptions such as CRS or similar quantum assumptions could evade these results. The case of composable security for the verification of multipartite entanglement is more positive; in Yehia et al. (2021) composable security was proven against dishonest preparation for GHZ states; while in Colisson et al. (2024) this was strengthened by proving composable security for verification of general graph states and allowing malicious parties to collude with the dishonest preparation.
@@ X.3 Single-photon sources
Despite presenting an inherent trade-off between high brightness and low multiphoton noise, Poisson-distributed sources are sufficient for demonstrations of two-partite quantum primitives in somewhat realistic conditions Pappa et al. (2014); Erven et al. (2014); Ng et al. (2012); Schiansky et al. (2023); Cao et al. (2024); Goorden et al. (2014); Collins et al. (2017); Clarke et al. (2012). On the other hand, it was recently shown that quantum dot single-photon sources (QDS), which in principle do not exhibit this trade-off, can provide performance enhancements for several protocols, mainly in terms of achievable distance Bozzio et al. (2022). This is partly thanks to their high single-photon purity and tunable coherence in the photon-number basis Loredo et al. (2019); Karli et al. (2024). Unlike for Poisson-distributed sources, the best performing QDS emit in the infrared Senellart et al. (2017), thus requiring frequency conversion to telecom wavelengths. On the other hand, QDS emitting directly in the telecom C-band are under active development Yu et al. (2023); Vyvlecka et al. (2023); Yang et al. (2024).
It is worth noting the existence of other single-photon sources such as organic molecules emitting at room temperature Murtaza et al. (2023); Toninelli et al. (2021) and defect centers in diamond Leifgen et al. (2014); Knall et al. (2022). To the best of our knowledge, their role in primitives beyond QKD has not yet been investigated.

### v2

@@ IX.2 Verifiable computation
Figure 8: Verifiable blind quantum computing. First proposed in Fitzsimons and Kashefi (2017) this figure was taken from the talk prepared for presenting Gheorghiu et al. (2019), and reproduced with permission from the authors. As defined in Eq. (17), the client generates qubits in a random basis string dictated by $\theta$. Red qubits denote the computation qubits, yellow qubits are dummy qubits that disentangle their neighbors from the graph and orange qubits are trap qubits. Trap qubits produce determinstic measurement outcomes, so the client can check if the received outcome $b_{i}$ matches the honest execution. For each qubit $i$, the client orders the server to perform measurement $M(\delta_{i})$ in a basis with angle $\delta_{i}$, and to send the outcome $b_{i}$ back to them.
@@ IX.3 Computation with classical clients
Assuming two quantum servers that share EPR pairs but cannot communicate with one another, a blind implementation of Shor’s factoring algorithm was performed for the number $N=15$, instructed by a fully classical client Huang et al. (2017). This involves the client splitting the computation into two parts, delegating one part to each of the photonic servers. Note that the CNOT gate is implemented differently for each of the two photonic servers, whose roles in the computation are asymmetric: for one server, it suffices to probabilistically post-select on the outputs of a polarizing beamsplitter, while the other server follows an implementation closer to Kiesel et al. (2005). By randomly alternating between computation and dummy tasks, the classical client is able to verify the behaviour of the servers and performing a CHSH test based on the protocol from Reichardt et al. (2013). Following up on the proposal from Huang et al. (2017), an extension to the blind factorization of $N=21$ was proposed Das and Sanders (2022). More recently, a simplified version of the classical client scheme from Mahadev (2018b) was demonstrated using a small trapped-ion quantum processor Stricker et al. (2024).
@@ X.1 Composability
Composable quantum cryptography is a large field and goes beyond the scope of this review [^16] Even the framework that needs to be used for composable security, suitable for quantum multiparty protocols, is an open area of research.. Most multiparty protocols cannot be composably secure in the plain model (without assumptions such as the CRS), and this persists in the quantum case. For instance, it was shown that weak coin flipping is not composably secure in the plain model Wu et al. (2025), while stand-alone sequential composition is proven secure. The case of relativistic quantum protocols (for bit commitment and coin flipping) were also shown to be non-composable in the plain model Vilasini et al. (2019). Introducing assumptions such as CRS or similar quantum assumptions could evade these results. The case of composable security for the verification of multipartite entanglement is more positive; in Yehia et al. (2021) composable security was proven against dishonest preparation for GHZ states; while in Colisson et al. (2024) this was strengthened by proving composable security for verification of general graph states and allowing malicious parties to collude with the dishonest preparation. However, limits on existing protocols were shown to be inevitable, in that quantum state verification cannot be both efficient and secure Wiesner et al. (2024).
@@ X.3 Single-photon sources
Despite presenting an inherent trade-off between high brightness and low multiphoton noise, Poisson-distributed sources are sufficient for demonstrations of two-partite quantum primitives in somewhat realistic conditions Pappa et al. (2014); Erven et al. (2014); Ng et al. (2012); Schiansky et al. (2023); Cao et al. (2024); Goorden et al. (2014); Collins et al. (2017); Clarke et al. (2012). On the other hand, it was recently shown that quantum dot single-photon sources (QDs), which in principle do not exhibit this trade-off, can provide performance enhancements for experimental QKD Zhang et al. (2025); Yang et al. (2024); Morrison et al. (2023) and mistrustful quantum primitives Bozzio et al. (2022); Vajner et al. (2024). This is partly thanks to their high single-photon purity, but also thanks to their tunable coherence in the photon-number basis Loredo et al. (2019); Karli et al. (2024); Hagen et al. (2025), which allows to circumvent the need for active phase randomization. Unlike for Poisson-distributed sources, the best performing QDs emit in the infrared Senellart et al. (2017), thus requiring frequency conversion to telecom wavelengths. On the other hand, QDs emitting directly in the telecom C-band are under active development Yu et al. (2023); Vyvlecka et al. (2023); Yang et al. (2024); Holewa et al. (2024).
It is worth noting the existence of other single-photon sources such as organic molecules emitting at room temperature Murtaza et al. (2023); Toninelli et al. (2021), defect centers in diamond Leifgen et al. (2014); Knall et al. (2022) and atomically-thin single-photon sources Gao et al. (2023). To the best of our knowledge, their role in primitives beyond QKD has not yet been investigated.